METHYL TORSIONAL LEVELS IN 9-METHYLANTHRACENE

\underline{MASAAKI BABA}, YUKI NOMA, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; SHUNJI KASAHARA, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; TAKAYA YAMANAKA, Institute for Molecular Science, Okazaki 444-8585, Japan; JON T. HOUGEN, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441, USA.

We have analysed energy levels of CH_{3} torsion in the S_{0} and S_{1} states of 9-methylanthracene (9MA) by high-resolution fluorescence excitation spectrum in a supersonic jet. 9MA has the G_{12} molecular symmetry, which is the same as toluene. The sixfold torsional barrier in toluene is very small, ^{a} and the barrier height has been accurately determined as V_{6}(S_{0}) = 4.874 cm^{−1} and V_{6}(S_{1}) = 26.376 cm^{−1} by microwave and ultrahigh-resolution laser spectroscopy. ^{b}^{,}^{c} In 9MA, we estimated the barrier height as V_{6}(S_{0}) ≈ 100 cm^{−1} and V_{6}(S_{1}) ≈ 50 cm^{−1}, which is remarkably larger than that in toluene. The difference of transition energies between the 0a′_{1} → 0a′_{1} and 1e′′ → 1e′′ is expected to be about 1 cm^{−1}. We have observed rotationally resolved ultrahigh-resolution spectrum of 9MA in a collimated molecular beam and confirmed this splitting. All of the observed rotational lines exhibit B-type selection rule, and the direction of transition moment has been shown to be parallel to the C_{3} axis of the CH_{3} group.

aP. J. Breen, J. A. Warren, E. R. Bernstein, and J. I. Seeman, J. Chem. Phys., 87, 1917 (1987)

bW. A. Kreiner, H. D. Rudolph, and B. T. Tan, J. Mol. Spectrosc., 48, 86 (1973)

cD. R. Borst and D. W. Pratt, J. Chem. Phys., 113, 3658 (2000)